OpenStudy (anonymous):

simplify x-7/x^3-127

OpenStudy (anonymous):

I'm trying to factor x^3-127...but am not sure if that is the right way to solve it

OpenStudy (anonymous):



OpenStudy (anonymous):

that's your question right

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

\(x^3\) is a cube. what about \(127\)?

OpenStudy (anonymous):

just 127

OpenStudy (whpalmer4):

If \(127=b^3, a^3=x^3\), then you have a difference of cubes: \[(a^3-b^3) = (a-b)(a^2+ab+b^2)\]However, 127 isn't a perfect cube.

OpenStudy (whpalmer4):

\[x^3-125 = (x-5)(x^2+5x+25)\] But that's not our problem. \[x^3-127\]is irreducible. Can't be factored.

OpenStudy (anonymous):

oops 125

OpenStudy (anonymous):

sorry

OpenStudy (whpalmer4):

Well, I just happen to have a writeup of that problem here somewhere…now where did I leave that? :-)

OpenStudy (anonymous):

Its here...I think I just saw it

OpenStudy (anonymous):

(x-5)(x^2+5x+25)

OpenStudy (whpalmer4):

We can test it out. Let x = 6, for example: \[(6)^3 - 125 = 216-125 = 91\]\[(6-5)(6^2+5(6)+25) = 1(36+30+25) = 91\checkmark \]

OpenStudy (anonymous):

so what will happen to x-7